# The Impact of Demonstrations on Multilingual In-Context Learning: A Multidimensional Analysis

### v1

@@ Abstract
In-context learning is a popular inference strategy where large language models solve a task using only a few labelled demonstrations without needing any parameter updates. Compared to work on monolingual (English) in-context learning, multilingual in-context learning is under-explored, and we lack an in-depth understanding of the role of demonstrations in this context. To address this gap, we conduct a multidimensional analysis of multilingual in-context learning, experimenting with 5 models from different model families, 9 datasets covering classification and generation tasks, and 56 typologically diverse languages. Our results reveal that the effectiveness of demonstrations varies significantly across models, tasks, and languages. We also find that Llama 2-Chat, GPT-3.5, and GPT-4 are largely insensitive to the quality of demonstrations. Instead, a carefully crafted template often eliminates the benefits of demonstrations for some tasks and languages altogether. These findings show that the importance of demonstrations might be overestimated. Our work highlights the need for granular evaluation across multiple axes towards a better understanding of in-context learning.
The Impact of Demonstrations on Multilingual In-Context
Learning: A Multidimensional Analysis
Miaoran Zhang1   Vagrant Gautam1   Mingyang Wang2,3   Jesujoba O. Alabi1 Xiaoyu Shen4Dietrich Klakow1Marius Mosbach1 1Saarland University, 2Bosch Center for AI, 3LMU Munich 4Eastern Institute of Technology, Ningbo mzhang@lsv.uni-saarland.de
@@ 1 Introduction
However, most recent work on in-context learning predominantly focuses on English, and the exploration of multilingual in-context learning generally lags behind. This is problematic, as results that apply to English might not hold for other languages, especially those which are less represented in LLM training data. While there have been a few studies on in-context learning that go beyond English, they either focus on benchmarking LLMs on multilingual tasks without in-depth exploration, e.g., MEGA (Ahuja et al., 2023) and BUFFET (Asai et al., 2023), or zoom in on specific capabilities such as machine translation (Zhu et al., 2023), code-switching (Zhang et al., 2023) or cross-lingual transfer (Shi et al., 2022; Tanwar et al., 2023; Agrawal et al., 2023).
In this work, we take a multidimensional approach (Ruder et al., 2022) that unifies these strands of research and comprehensively evaluate the multilingual in-context learning abilities of LLMs. We focus on dissecting the actual impact of demonstrations across various models, tasks and languages. Our research questions are:
@@
Figure 1: An overview of the components of in-context learning (§2). Sources of variation include tasks, languages, models, and the template, i.e., the task instruction, patterns for formatting inputs, and verbalized labels.
We address our research questions by evaluating 5 LLMs including base models that are only pre-trained on unlabelled text corpora (XGLM and Llama 2), and chat models that are further refined with instruction tuning and reinforcement learning (Llama 2-Chat, GPT-3.5, and GPT-4). We evaluate on 9 multilingual datasets, covering 56 typologically different languages.
Our main findings are: (1) The effectiveness of demonstrations varies widely depending on the model, task, and language used. For base models, in-context learning barely outperforms zero-shot learning on many tasks. In general, in-context learning matters more for generation tasks with loosely-specified prompts; (2) Even with sophisticated demonstration selection methods, in-context learning is not always beneficial and can sometimes be worse than using no demonstrations at all; (3) Chat models are less sensitive to seeing correctly-labelled demonstrations than base models, suggesting that for the former, demonstrations primarily help the model understand the task format, while for the latter, demonstrations also impart task-specific knowledge; (4) Using a formatting-focused template can even eliminate the need for demonstrations with chat models.
In sum, we suggest that the benefits of adding demonstrations may be overestimated. Future work on in-context learning should carefully compare their results with zero-shot learning and on multiple templates, languages and models to more comprehensively understand its performance.
2 In-context learning
In-context learning (ICL) is a popular inference strategy where models solve [^1] The extent to which models actually “solve” tasks is an open question as ICL, similar to fine-tuning, has generalization issues despite its impressive results (Mosbach et al., 2023). Regardless, we use the word “solve” in the rest of this paper for simplicity. a task without any parameter updates (Brown et al., 2020). Instead, the model performs the task by conditioning on labelled demonstrations. Demonstrations are typically formatted using “pattern-verbalizer pairs,” as this has been shown to be effective in eliciting good task performance (Schick and Schütze, 2021; Bach et al., 2022). Here, a pattern is used to format the input for the model, and a verbalizer maps the label to a textual representation. Additionally for instruction-tuned LLMs, a task instruction is often added to provide information about the task beyond individual demonstrations (Mishra et al., 2022b; Wang et al., 2022; Ouyang et al., 2022).
@@ 2 In-context learning
Taken together, the pattern, the verbalizer and the optional task instruction comprise the template with which demonstrations as formatted as the input prompt for the model during in-context-learning. The effectiveness of demonstrations is thus linked with the template used to present them to the model.
Figure 2: Average performance across languages for classification tasks with different numbers of demonstrations. We average and report standard deviations over 3 seeds for all models except GPT-4. Note that the standard deviations are relatively small, possibly because of averaging over languages.
@@ 3 Experimental setup
We evaluate two types of LLMs: pre-trained base models and chat models. Our base models are XGLM (Lin et al., 2022) and Llama 2 (Touvron et al., 2023). Our chat models are Llama 2-Chat, GPT-3.5 (Ouyang et al., 2022) and GPT-4 (OpenAI et al., 2023). Specifically, we use xglm-7.5B, Llama-2-13b, and Llama-2-13b-chat on Huggingface (Wolf et al., 2020), and we access gpt-3.5-turbo-16k and gpt-4-32k APIs via Microsoft Azure. [^2] We also experiment with BLOOMZ and mT0 (Muennighoff et al., 2023). Results in Appendix B.1 show that their zero-shot performance significantly surpasses few-shot performance, which we ascribe to their training scheme.
@@
For each test example, we select $k\in\{0,2,4,8\}$ [^3] For QA datasets, we select a maximum of $4$ demonstrations due to context size limitations. different demonstrations, which are randomly sampled unless otherwise specified. All demonstrations are in the same language as the test example (Shi et al., 2023b), and all templates are in English, as English templates have been shown to perform better than native language templates (Lin et al., 2022; Ahuja et al., 2023). We employ appropriate task-specific templates for different model types. All templates and data splits are shown in Appendix A.2.
Figure 3: Average performance across languages for generation tasks with different numbers of demonstrations. We average and report standard deviations over 3 seeds for all models except GPT-4. Note that the standard deviations are relatively small, possibly because of averaging over languages. en-xx: translating from English to a target language, xx-en: translating from a target language to English.
@@
For classification tasks, we report the rank classification accuracy [^4] The scoring function is the average of per-token log probabilities (ignoring the common prefix of different candidates). The candidate with highest score is chosen as the prediction. for open-source base models (Muennighoff et al., 2023; Lin et al., 2022). For chat models, we measure the exact match between generated outputs [^5] We extract verbalized labels from the generated outputs using regular expressions before calculating the exact match. and verbalized labels (Ahuja et al., 2023). As for generation tasks, we use the F1 score for QA datasets, and ChrF++ score (Popović, 2017) for MAFAND. Implementation details for our evaluation are provided in Appendix A.3.
@@ 4 Do (more) demonstrations benefit multilingual performance?
In this section, we systematically compare ICL and zero-shot learning as this question is underexplored in previous studies of multilingual ICL (Ahuja et al., 2023; Asai et al., 2023). We examine model performance on diverse multilingual tasks while varying the number of demonstrations, and show the results for classification tasks in Figure 2 and generation tasks in Figure 3.
We begin with the overall trends across models and datasets. OpenAI’s GPT-3.5 and GPT-4 models achieve the best multilingual in-context learning performance on all our datasets, which is unsurprising as they are currently the state-of-the-art on a large suite of NLP benchmarks. The next best models are Llama 2 and Llama 2-Chat, which demonstrate comparable or superior performance to the multilingual XGLM model despite being trained primarily on English corpora (Touvron et al., 2023). This suggests that their task-solving abilities can transfer across languages.
An important pattern across datasets and models is that in-context learning does not always improve over zero-shot learning – in particular, it helps with generation tasks, but results on classification tasks are mixed. For the AfriSenti dataset, many models show large improvements with ICL. However, with other tasks such as PAWS-X, XNLI and IndicXNLI, the same models perform drastically worse compared to the zero-shot setting. We also see marginal improvements in some cases, e.g., XGLM and Llama 2 on XCOPA. In comparison to chat models, the addition of demonstrations often reduces the performance of base models across many tasks. When examining the cases where ICL improves performance, we see that improvements saturate quickly with 2 to 4 demonstrations. This aligns with Chen et al. (2023), who found that reducing the number of demonstrations to one does not significantly deteriorate chain-of-thought reasoning.
Table 1: Performance difference of 4-shot ICL with top-k vs. random selection. Positive numbers show that top-k is better than random (expected), and highlighted cells show where top-k is even worse than random selection. $\triangledown$: top-k performance is worse than zero-shot learning. For random, we average over 3 seeds (except for GPT-4). We cannot calculate significance as top-k selects a single set of examples per test sample.
@@
Table 2: Performance difference of 4-shot ICL with random vs. random-corrupted selection. Positive numbers show that random is better than random-corrupted (expected), and highlighted cells show where corrupted labels perform even better than ground-truth labels. We average over 3 seeds (except for GPT-4). ***: a significant difference ($p=0.05$).
@@
Looking at the improvements over zero-shot performance (for all models and languages combined) across tasks in Figure 4, we observe that there are large fluctuations between individual languages that are not captured by the average. The PAWS-X dataset in particular shows an average degradation, but in fact some languages benefit from ICL while others degrade. For a more nuanced understanding of language-specific differences within a task, we zoom into this dataset in Figure 5 to inspect these language-specific differences. [^6] Additional plots for other datasets are in Appendix B.2. We see that languages and models can behave very differently even on just one dataset, and a pattern which holds for one language with one model does not necessarily apply to a different language. For example, on Japanese and Korean, the ICL performance of Llama 2 is slightly higher than its zero-shot performance, but for other languages, demonstrations degrade performance. In sum, the effectiveness of demonstrations varies widely depending on the model, task, and language.

### v2

@@ Abstract
In-context learning is a popular inference strategy where large language models solve a task using only a few labeled demonstrations without needing any parameter updates. Although there have been extensive studies on English in-context learning, multilingual in-context learning remains under-explored, and we lack an in-depth understanding of the role of demonstrations in this context. To address this gap, we conduct a multidimensional analysis of multilingual in-context learning, experimenting with 5 models from different model families, 9 datasets covering classification and generation tasks, and 56 typologically diverse languages. Our results reveal that the effectiveness of demonstrations varies significantly across models, tasks, and languages. We also find that strong instruction-following models including Llama 2-Chat, GPT-3.5, and GPT-4 are largely insensitive to the quality of demonstrations. Instead, a carefully crafted template often eliminates the benefits of demonstrations for some tasks and languages altogether. These findings show that the importance of demonstrations might be overestimated. Our work highlights the need for granular evaluation across multiple axes towards a better understanding of in-context learning. [^1] We release our code publicly at https://github.com/uds-lsv/multilingual-icl-analysis.
The Impact of Demonstrations on Multilingual In-Context Learning:
A Multidimensional Analysis
Miaoran Zhang1   Vagrant Gautam1   Mingyang Wang2,3,4   Jesujoba O. Alabi1 Xiaoyu Shen5thanks:   Corresponding author. Dietrich Klakow1Marius Mosbach6 1Saarland University, Saarland Informatic Campus 2Bosch Center for AI 3LMU Munich 4Munich Center for Machine Learning (MCML) 5Eastern Institute of Technology, Ningbo   6Mila, McGill University {mzhang,vgautam,jalabi,dietrich.klakow}@lsv.uni-saarland.de mingyang@cis.lmu.de xyshen@eitech.edu.cn  marius.mosbach@mila.quebec
@@ 1 Introduction
However, most recent work on in-context learning predominantly focuses on English, and the exploration of multilingual in-context learning generally lags behind. This is problematic, as results that apply to English might not hold for other languages, especially those that are less represented in LLM training data. While there have been a few studies on in-context learning that go beyond English, they either focus on benchmarking LLMs on multilingual tasks without in-depth exploration, e.g., MEGA (Ahuja et al., 2023) and BUFFET (Asai et al., 2023), or zoom in on specific capabilities such as mathematical reasoning (Shi et al., 2023b), machine translation (Zhu et al., 2023; Agrawal et al., 2023), or code-switching (Zhang et al., 2023).
In this work, we take a multidimensional approach (Ruder et al., 2022) that unifies these strands of research and comprehensively evaluate the multilingual in-context learning abilities of LLMs. We focus on dissecting the actual impact of in-context demonstrations, which is crucial for understanding model behaviour. Our research covers various models, tasks, and languages, and we seek to answer the following research questions:
@@
Figure 1: An overview of the components of multilingual in-context learning (§2) with a comparison to zero-shot learning. Sources of variation include tasks, languages, models, and the template, i.e., the task instruction, patterns for formatting inputs, and verbalized labels.
Specifically, we address our research questions by evaluating 5 LLMs including base models that are only pre-trained on unlabeled text corpora (XGLM and Llama 2), and chat models that are further refined with instruction tuning and reinforcement learning (Llama 2-Chat, GPT-3.5, and GPT-4). We evaluate on 9 multilingual datasets that include both classification and generation tasks, covering 56 typologically different languages.
Our main findings are: (1) The effectiveness of demonstrations varies widely depending on the model, task, and language used. For base models, in-context learning barely outperforms zero-shot learning on many tasks. In general, in-context learning matters more for generation tasks with loosely-specified prompts; (2) Even with sophisticated demonstration selection methods, in-context learning is not always beneficial and can sometimes be worse than using no demonstrations at all; (3) Chat models are less sensitive to seeing correctly-labeled demonstrations than base models, suggesting that for the former, demonstrations primarily help the model understand the task format, while for the latter, demonstrations also impart task-specific knowledge; (4) Using a formatting-focused template can even eliminate the need for demonstrations with chat models. The relative significance of demonstrations versus prompt templates varies based on inherent model capabilities.
In sum, we suggest that the benefits of adding demonstrations may be overestimated. Future work on in-context learning should carefully compare their results with zero-shot learning and on multiple templates to faithfully represent its effectiveness. Given the vast variance across models, tasks, and languages, it is also important to cautiously frame claims about in-context learning.
2 Preliminaries
2.1 In-context learning
In-context learning (ICL) is a popular inference strategy where models solve [^2] The extent to which models actually “solve” tasks is an open question as ICL, similar to fine-tuning, has generalization issues despite its impressive results (Mosbach et al., 2023). Regardless, we use the word “solve” in the rest of this paper for simplicity. a task without any parameter updates (Brown et al., 2020). Instead, the model performs the task by conditioning on labeled demonstrations. Demonstrations are typically formatted using “pattern-verbalizer pairs,” as this has been shown to be effective in eliciting good task performance (Schick and Schütze, 2021; Bach et al., 2022). Here, a pattern is used to format the input for the model, and a verbalizer maps the label to a textual representation. Additionally for instruction-tuned LLMs, a task instruction is often added to provide information about the task beyond individual demonstrations (Mishra et al., 2022b; Wang et al., 2022; Ouyang et al., 2022).
@@ 2.1 In-context learning
Taken together, the pattern, the verbalizer and the optional task instruction comprise the template with which demonstrations and the test sample are formatted as the input prompt for model inference. The effectiveness of demonstrations is thus linked with the template used to present them to the model.
2.2 Multilingual prompting
Previous studies highlight that the selection of demonstrations and prompt templates can significantly influence model performance (Liu et al., 2022; Fu et al., 2023b; Sclar et al., 2024). In multilingual in-context learning, the variation in input prompts is further complicated by the language of demonstrations, templates and test samples, all of which are important design choices.
For the template language, Lin et al. (2022) and Ahuja et al. (2023) found that English templates generally perform better than native language templates, possibly due to superior instruction-following abilities on existing LLMs on English compared to other languages. Following this, we use English templates in our study.
For the language of few-shot demonstrations and test samples, there are three popular settings. Given a test sample in a certain language, the most straightforward approach is to use demonstrations in the same language (referred to as in-language demonstrations). This setting directly measures the model’s inherent ability to solve problems in that language. Another choice is to use English demonstrations regardless of the language of the test sample. This is a cross-lingual transfer setup, where the goal is to transfer knowledge from a pivot language to a target language via in-context learning. As highlighted in Shi et al. (2023b) and Ahuja et al. (2023), in-language demonstrations often outperform English demonstrations on diverse multilingual tasks. Yet another option is to translate the test sample into English – an approach called translate-test, where the demonstrations are also in English. While translate-test leads to strong performance (Ahuja et al., 2023), this approach heavily relies on a translation system for data processing and centers the English proficiency of LLMs. In this work, we are interested in dissecting the intrinsic multilingual capabilities of LLMs, therefore we choose to use in-language demonstrations.
All these design choices are represented visually in Figure 1, which gives an overview of multilingual in-context learning. Detailed setup information is provided in the next section.
Figure 2: Average performance across languages with different numbers of demonstrations. We average and report standard deviations over 3 seeds for all models except GPT-4. Note that the standard deviations are relatively small, possibly because of averaging over languages. en-xx: translating from English to another language, xx-en: translating from another language to English.
@@ 3 Experimental setup
We evaluate two types of LLMs: pre-trained base models and chat models. Our base models include XGLM (Lin et al., 2022) and Llama 2 (Touvron et al., 2023). Our chat models are Llama 2-Chat, GPT-3.5 (Ouyang et al., 2022) and GPT-4 (OpenAI et al., 2023). Specifically, we use xglm-7.5B, Llama-2-13b, and Llama-2-13b-chat on Huggingface (Wolf et al., 2020), and we access gpt-3.5-turbo-16k and gpt-4-32k APIs via Microsoft Azure. [^3] We also experiment with BLOOMZ and mT0 (Muennighoff et al., 2023). Results in Appendix B.1 show that their zero-shot performance significantly surpasses few-shot performance, which we ascribe to their training scheme.
@@
For each test sample, we select $k\in\{0,2,4,8\}$ [^4] For QA datasets, we select a maximum of $4$ demonstrations due to context size limitations. different demonstrations, which are randomly sampled unless otherwise specified. All demonstrations are in the same language as the test sample, and all templates are in English. We employ appropriate task-specific templates for different model types. All templates and data splits are shown in Appendix A.2.
@@
For classification tasks, we report the rank classification accuracy [^5] The scoring function is the average of per-token log probabilities (ignoring the common prefix of different candidates). The candidate with the highest score is chosen as the prediction. for open-source base models (Muennighoff et al., 2023; Lin et al., 2022). For chat models, we measure the exact match between generated outputs [^6] We extract verbalized labels from the generated outputs using regular expressions before calculating the exact match. and verbalized labels (Ahuja et al., 2023). As for generation tasks, we use the F1 score for QA datasets and ChrF++ score (Popović, 2017) for MAFAND. Implementation details for our evaluation are provided in Appendix A.3.
@@ 4 Do (more) demonstrations benefit multilingual performance?
In this section, we systematically compare ICL and zero-shot learning as this question is under-explored in previous studies of multilingual ICL (Ahuja et al., 2023; Asai et al., 2023). We examine model performance on diverse multilingual tasks while varying the number of demonstrations, and show the results for classification tasks and generation tasks in Figure 2.
We begin with the overall trends across models and datasets. OpenAI’s GPT-3.5 and GPT-4 models achieve the best multilingual in-context learning performance on all our datasets, which is unsurprising as they are currently the state-of-the-art on a large suite of NLP benchmarks. The next best models are Llama 2 and Llama 2-Chat, which demonstrate comparable or superior performance to the multilingual XGLM model despite being trained primarily on English corpora (Touvron et al., 2023). This indicates that their task-solving abilities can transfer across languages. Regardless of the model, however, performance on the AfriSenti and MAFAND datasets, particularly when translating English to African languages, lags significantly behind other tasks, showing that language discrepancies remain even in the best models.
@@
Table 1: Performance difference of 4-shot ICL with top-k vs. random selection. Positive numbers show that top-k is better than random (expected), and highlighted cells show where top-k is even worse than random selection. $\triangledown$: top-k performance is even worse than zero-shot learning. For random, we average over 3 seeds (except for GPT-4).